A train leaves a station at 6am and covers a distance of 300km by 12noon,what is its average speed for the journey?
Answer by checkley77(12844)

12-6=6 HOURS FOR THE TIME OF THE TRAINS JOURNEY.
D=RT
300=R*6
R=300/6
R=50 MPH. ANS.